# DMVC-Tracker: Distributed Multi-Agent Trajectory Planning for Target Tracking Using Dynamic Buffered Voronoi and Inter-Visibility Cells

### v1

@@ IV-B2 Acute case
$1\leq\alpha_{a}^{ij}\leq\min\Bigg{(}\frac{\bigg{\|}\det\bigg{(}\begin{bmatrix}(\textbf{x}_{c0}^{i}-\textbf{x}_{q0})^{\top}\\ (\textbf{x}_{c0}^{j}-\textbf{x}_{q0})^{\top}\end{bmatrix}\bigg{)}\bigg{\|}}{r_{c}\max(\|\textbf{x}_{c0}^{i}-\textbf{x}_{q0}\|,\|\textbf{x}_{c0}^{j}-\textbf{x}_{q0}\|)},$ (17)
$\quad\quad\quad\quad\quad\quad\sqrt{2\bigg{(}1+\frac{(\textbf{x}_{c0}^{i}-\textbf{x}_{q0})^{\top}(\textbf{x}_{c0}^{j}-\textbf{x}_{q0})}{\|\textbf{x}_{c0}^{i}-\textbf{x}_{q0}\|\|\textbf{x}_{c0}^{j}-\textbf{x}_{q0}\|}\bigg{)}}\ \Bigg{)},$
@@
$\mathcal{H}^{ij}_{a1}(t)=$$\bigg{\{}\textbf{x}(t)\in\mathbb{R}^{2}|z^{ij}\det\bigg{(}\begin{bmatrix}(\textbf{x}(t)-\textbf{x}_{q}(t))^{\top}\\ (\textbf{x}_{c0}^{j}-\textbf{x}_{q0})^{\top}\end{bmatrix}\bigg{)}+$
$\quad\quad\quad\alpha_{a}^{ij}r_{c}\|\textbf{x}_{c0}^{j}-\textbf{x}_{q0}\|\leq 0\bigg{\}},$ (18a)
@@
$\ \textbf{x}_{q}(t)-\alpha_{a}^{ij}r_{c}\frac{\|\textbf{x}_{c0}^{j}-\textbf{x}_{q0}\|(\textbf{x}_{c0}^{i}-\textbf{x}_{q0})}{\bigg{\|}\det\Big{(}\begin{bmatrix}(\textbf{x}_{c0}^{i}-\textbf{x}_{q0})^{\top}\\ (\textbf{x}_{c0}^{j}-\textbf{x}_{q0})^{\top}\end{bmatrix}\Big{)}\bigg{\|}}\bigg{)}\leq 0\bigg{\}}$ (18b)
@@
$\sin(\theta_{a}^{ij})=\frac{\bigg{\|}\det\Big{(}\begin{bmatrix}(\textbf{x}_{c0}^{i}-\textbf{x}_{q0})^{\top}\\ (\textbf{x}_{c0}^{j}-\textbf{x}_{q0})^{\top}\end{bmatrix}\Big{)}\bigg{\|}}{\alpha_{a}^{ij}\|\textbf{x}_{c0}^{i}-\textbf{x}_{q0}\|\|\textbf{x}_{c0}^{j}-\textbf{x}_{q0}\|},$ (19)
@@ V Tracking Trajectory Planning
In this section, we formulate the tracking trajectory generation using Bernstein polynomial motion primitives. First, we sample a bundle of motion primitives and filter out the primitives that do not satisfy constraints. Then, the best trajectory is selected.
@@ V-B1 Collision Check
$\mathcal{S}_{m}^{i}$ is the $m$-th corridors generated by [10], $\textbf{C}_{m}^{i}$ is the $m$-th Bernstein coefficients, which is split from $\textbf{C}^{i}$. $\Xi$ is a set of points in Bernstein coefficients. Please refer to [10] for the details. Similarly to (22a), we verify whether the control points of (22d) belong to the affine spaces by using the convex hull property to determine if trajectories are located within regions where inter-collision does not occur. The primitives that pass the tests in (22) satisfy (3) and (7).
@@ V-B2 Visibility Check
To enable the $i$-th agent to see the target, we check whether the primitives satisfy (23a), (23b), and (23c) to avoid occlusion by static obstacles, dynamic obstacles, and the other agents, respectively.
@@
The primitives that pass the checks in (23) satisfy (4) and (8). (23b) represents a condition that Line-of-Sight does not collide with dynamic obstacles, and the inequality in (23b) can be reformulated as (24a).
$\epsilon^{2}\sigma_{1}(t)+2\epsilon(1-\epsilon)\sigma_{2}(t)+(1-\epsilon)^{2}\sigma_{3}(t)\geq 0,$ (24a)
$\sigma_{1}(t)=\|\textbf{x}_{c}(t)-\textbf{x}_{o}^{k}(t)\|^{2}-(r_{o}+r_{c})^{2},$ (24b)
$\sigma_{2}(t)=(\textbf{x}_{c}(t)-\textbf{x}_{o}^{k}(t))^{\top}(\textbf{x}_{q}(t)-\textbf{x}_{o}^{k}(t))$ (24c)
$\quad\quad\quad+(r_{o}+\min(r_{q},r_{c}))^{2}-2r_{o}^{2},$
$\sigma_{3}(t)=\|\textbf{x}_{q}(t)-\textbf{x}_{o}^{k}(t)\|^{2}-(r_{o}+r_{q})^{2}$ (24d)
Since $\epsilon^{2}$, $2\epsilon(1-\epsilon)$, and $(1-\epsilon)^{2}$ are nonnegative for $\forall\epsilon\in[0,1]$, if $\sigma_{1}(t)$, $\sigma_{2}(t)$, $\sigma_{3}(t)\geq 0$ for $\forall t\in[0,T]$, (24a) holds. Fig. 3 shows that the proposed check method is less conservative compared to our previous work [10]. By employing the less conservative feasibility check, our planner can find more feasible motions.
@@ V-B3 Distance Check
$\|\textbf{x}_{c}^{i}(t)-\textbf{x}_{q}(t)\|^{2}-d_{\min}^{2}\geq 0,$ (25a)
$d_{\max}^{2}-\|\textbf{x}_{c}^{i}(t)-\textbf{x}_{q}(t)\|^{2}\geq 0$ (25b)
@@
Scenarios Environments Planner Safety Metrics Visibility Metrics
$\chi_{1}(t)$ (28a) [m] $\chi_{2}(t)$ (28b) [m] $\chi_{3}(t)$ (28c) [m] $\phi_{1}(t)$ (28d) [m] $\phi_{2}(t)$ (28e) [m]
@@ V-B4 Dynamical Limit Check
To ensure the drone does not exceed dynamic limits (6), we verify whether the primitives satisfy the following inequalities. (26c) is calculated under the assumption that the agents directly head toward the target.
$v_{\max}^{2}-\|\dot{\textbf{x}}_{c}^{i}(t)\|^{2}\geq 0,$ (26a)
$a_{\max}^{2}-\|\ddot{\textbf{x}}_{c}^{i}(t)\|^{2}\geq 0,$ (26b)
$\dot{\psi}_{c}(t)=\frac{\det\Big{(}\begin{bmatrix}(\textbf{x}_{q}(t)-\textbf{x}_{c}^{i}(t))^{\top}\\ (\dot{\textbf{x}}_{q}(t)-\dot{\textbf{x}}_{c}^{i}(t))^{\top}\end{bmatrix}\Big{)}}{\|\textbf{x}_{q}(t)-\textbf{x}_{c}^{i}(t)\|^{2}},$ (26c)
$-\dot{\psi}_{\max}\leq\dot{\psi}_{c}(t)\leq\dot{\psi}_{\max}$ (26d)
@@ VI Validations
In this section, the proposed method is validated through various target-tracking settings. We measure the distance between the drone and environments, defined as (28a), the distance among agents (28b), and the distance between the target and drone, defined as (28c), to evaluate drone safety. Also, we measure the distance between the Lines-of-Sight and environments (28d) and the distance between the Lines-of-Sight and the other agents (28e) to assess the target visibility.
@@
$\chi_{3}(t)=\min_{i\in\mathcal{I}_{c}}\min_{\begin{subarray}{c}\textbf{x}(t)\in\mathcal{C}_{i}(t)\\ \textbf{y}(t)\in\mathcal{T}(t)\end{subarray}}\|\textbf{x}(t)-\textbf{y}(t)\|,$ (28c)
$\phi_{1}(t)=\min_{i\in\mathcal{I}_{c}}\min_{\begin{subarray}{c}\textbf{x}(t)\in\mathcal{L}(\textbf{x}_{c}^{i}(t),\textbf{x}_{q}(t))\\ \textbf{y}(t)\in\mathcal{O}(t)\end{subarray}}\|\textbf{x}(t)-\textbf{y}(t)\|,$ (28d)
$\phi_{2}(t)=\min_{i\in\mathcal{I}_{c}}\min_{\begin{subarray}{c}\textbf{x}(t)\in\mathcal{L}(\textbf{x}_{c}^{i}(t),\textbf{x}_{q}(t))\\ \textbf{y}(t)\in\mathcal{C}^{i}(t)\end{subarray}}\|\textbf{x}(t)-\textbf{y}(t)\|$ (28e)
Using the above performance metrics, we validate the operability of the proposed planner under two environmental conditions: 1) unstructured static obstacles, and 2) dynamic obstacles. Through simulations and hardware experiments, we show successful target tracking. Table II shows the reported performance, and the details of the tests are explained in the following subsections. In addition, we validate the effectiveness of the proposed method for multi-agent tracking missions through a comparative analysis.
@@ VI-A Simulations
Scenario 1 (Unstructured Environment): We compare the proposed planner with the state-of-the-art planner [21] in an environment with various shapes of static obstacles, as shown in Fig. 4. As Table II shows, the trackers controlled by the baseline [21] experience inter-agent occlusions and collisions with the target several times, despite finely tuned parameters to adapt to the narrow tracking environment. In contrast, our planner successfully tracks the target without collision and occlusion by thoroughly checking safety and visibility constraints, including the tracking distance $(d_{\min},d_{\max})$ and DIVC.
@@ VI-C Comparison Analysis
The performance of the proposed approach is investigated through two tests. First, we evaluate the effectiveness of the DBVC and the DIVC. Second, we compare the performance of our approach with the planner presented in [10]. In each test, we define success as the absence of occlusion or collision until all objects come to a stop. We conduct 1000 tests for each tracking setup and measure the success rate.
VI-C1 Ablation Study
We validate the effectiveness of the DBVC and the DIVC by comparing the success rates with cases where these two cells are not applied and where static versions of the cells are used. Specifically, when the cells are not used, each tracker treats the neighboring trackers as dynamic obstacles. In the case of using static cells, we use $\mathcal{H}_{s}^{ij}$, $\mathcal{H}_{\mu 1}^{ij}$, and $\mathcal{H}_{\mu 2}^{ij},\mu=o$ or $a$, instead of $\mathcal{H}_{s}^{ij}(t)$, $\mathcal{H}_{\mu 1}^{ij}(t)$, and $\mathcal{H}_{\mu 2}^{ij}(t)$.
@@ VI-C1 Ablation Study
VI-C2 Benchmark Test
We conduct a benchmark test to validate the superiority of our planner in a dynamic obstacle environment. We compare our method with a noncooperative approach and a conservative approach. For the noncooperative approach, we use the BPMP-Tracker [10]. Since the method in [10] is designed for single-agent tracking, we adapt the problem setting so that each tracker treats neighboring agents as obstacles. The conservative approach utilizes the DBVC and the DIVC but applies the conservative feasibility check methods from [10].
@@ VII Conclusion
We presented a distributed multi-agent trajectory generation method for aerial tracking, which prevents not only occlusion and collision caused by obstacles but also inter-occlusion and inter-collision among agents. Dynamic Buffered Voronoi Cells (DBVC) and Dynamic Inter-Visibility Cells (DIVC) were constructed by dividing the space into multiple cells to avoid inter-agent collision and inter-agent occlusion, respectively. Since both cells are designed based on the shared current positions of agents and the predicted target’s trajectory, it enables distributed planning. We achieved fast computation by combining the DBVC and the DIVC with the Bernstein-polynomial-motion-primitive-based trajectory planner and validated the operability of our planner through tests under various tracking conditions. Lastly, we confirmed that the proposed method achieved a higher success rate in tracking missions in environments with dynamic obstacles, outperforming the state-of-the-art method in noncooperative approaches and methods employing conservative feasibility checks.

### v2

@@ IV-B2 Acute case
$1\leq\alpha_{a}^{ij}\leq\min\bigg{(}\frac{\big{\|}\det([(\textbf{x}_{cq0}^{i})^{\top};(\textbf{x}_{cq0}^{j})^{\top}])\big{\|}}{r_{c}\max(\|\textbf{x}_{cq0}^{i}\|,\|\textbf{x}_{cq0}^{j}\|)},$ (14)
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\sqrt{2(1+(\textbf{n}_{cq0}^{i})^{\top}\textbf{n}_{cq0}^{j})}\bigg{)}$
@@
$\mathcal{H}^{ij}_{a1}(t)=$$\big{\{}\textbf{x}(t)\in\mathbb{R}^{2}|z^{ij}\det\big{(}[(\textbf{x}(t)-\textbf{x}_{q}(t))^{\top};(\textbf{x}_{cq0}^{j})^{\top}]\big{)}+$
$\quad\quad\quad\alpha_{a}^{ij}r_{c}\|\textbf{x}_{cq0}^{j}\|\leq 0\big{\}},$ (15a)
@@
$\ \textbf{x}_{q}(t)-\frac{\alpha_{a}^{ij}r_{c}\textbf{n}_{cq0}^{i}}{\big{\|}\det\big{(}[(\textbf{n}_{cq0}^{i})^{\top};(\textbf{n}_{cq0}^{j})^{\top}\big{)}\big{\|}}\big{)}\leq 0\bigg{\}}$ (15b)
@@
$\sin(\theta_{a}^{ij})=\frac{1}{\alpha_{a}^{ij}}\big{\|}\det\big{(}[(\textbf{n}_{cq0}^{i})^{\top};(\textbf{n}_{cq0}^{j})^{\top}]\big{)}\big{\|},$ (16)
@@ V Tracking Trajectory Planning
In this section, we formulate the tracking trajectory generation using Bernstein polynomial motion primitives. We begin by sampling a set of motion primitives and filtering out those that do not satisfy the given constraints. Next, we select the optimal trajectory. To enhance computational efficiency during both the feasibility check and trajectory selection processes, we fully exploit the convex hull property and the integral property of the Bernstein polynomial, respectively.
@@ V-B1 Collision Check
$\mathcal{S}_{m}^{i}$ is the $m$-th visible and safe corridors generated by [8], $\textbf{C}_{m}^{i}$ is the $m$-th Bernstein coefficients, which is split from $\textbf{C}^{i}$. $\Xi$ is a set of points in Bernstein coefficients. For (19a), we utilize convex hull property, and since the left-hand-side of (19b) and (19c) can be expressed as Bernstein polynomials, we select primitives with nonnegative coefficients for all bases. Similarly to (19a), we verify whether the control points of (19d) belong to the intersections of the affine spaces by using the convex hull property to determine if trajectories are located within DBVC where inter-collision does not occur. The primitives that pass the tests in (19) satisfy (3).
@@ V-B2 Visibility Check
We check whether the primitives of the $i$-th agent satisfy (20a), (20b), and (20c) to avoid occlusion by static obstacles, dynamic obstacles, and the other agents, respectively, and consequently satisfy (4).
@@
(20a) is a necessary condition for (19a). The convex hull property is used to check (20c), the DIVC constraints. (20b) means that Line-of-Sight does not collide with dynamic obstacles, and the left-hand side of (20b) can be reformulated as follows.
[…]
$\quad\quad\quad\quad\quad\quad\geq\epsilon^{2}\sigma_{1}^{\prime}(t)+2\epsilon(1-\epsilon)\sigma_{2}^{\prime}(t)+(1-\epsilon)^{2}\sigma_{3}^{\prime}(t),$
$\text{where}\ \sigma_{1}(t)=\|\textbf{x}_{c}^{i}(t)-\textbf{x}_{o}^{k}(t)\|^{2}-r_{o}^{2},$
$\quad\quad\ \ \sigma_{2}(t)=(\textbf{x}_{c}(t)-\textbf{x}_{o}^{k}(t))^{\top}(\textbf{x}_{q}(t)-\textbf{x}_{o}^{k}(t))-r_{o}^{2},$
$\quad\quad\ \ \sigma_{3}(t)=\|\textbf{x}_{q}(t)-\textbf{x}_{o}^{k}(t)\|^{2}-r_{o}^{2},$
$\quad\quad\ \ \sigma_{1}^{\prime}(t)=\sigma_{1}(t)-(r_{o}+r_{c})^{2}+r_{o}^{2},$
$\quad\quad\ \ \sigma_{2}^{\prime}(t)=\sigma_{2}(t)+(r_{o}+\min(r_{q},r_{c}))^{2}-r_{o}^{2},$
$\quad\quad\ \ \sigma_{3}^{\prime}(t)=\sigma_{3}(t)-(r_{o}+r_{q})^{2}+r_{o}^{2}$
Since $\epsilon^{2}$, $2\epsilon(1-\epsilon)$, and $(1-\epsilon)^{2}$ are nonnegative for $\forall\epsilon\in[0,1]$, (20b) holds if the terms multiplied by them are also nonnegative. This paper examines whether $\sigma_{1}^{\prime}(t)$, $\sigma_{2}^{\prime}(t)$, $\sigma_{3}^{\prime}(t)\geq 0$ hold, whereas our previous work [8] checks whether $\sigma_{1}(t)$, $\sigma_{2}(t)$, $\sigma_{3}(t)\geq 0$ hold. By tightening the conditions for $\sigma_{1}(t)$ and $\sigma_{3}(t)$, the proposed work relaxes condition for $\sigma_{2}(t)$. Fig. 3 shows that the proposed check method is less conservative compared to [8]. By employing the less conservative feasibility check, our planner can find more feasible motions.
@@ V-B3 Distance Check
$\|\textbf{x}_{c}^{i}(t)-\textbf{x}_{q}(t)\|^{2}-d_{\min}^{2}\geq 0,\ d_{\max}^{2}-\|\textbf{x}_{c}^{i}(t)-\textbf{x}_{q}(t)\|^{2}\geq 0$ (22)
@@
Scenarios Environments Planner Safety Metrics [m] Visibility Metrics [m]
Obstacle (25a) Agent (25b) Target (25c) Obstacle (26a) Agent (26b)
@@ V-B4 Dynamical Limit Check
To ensure the drone does not exceed dynamic limits (6), we verify whether the primitives satisfy the following inequalities. (23b) is calculated under the assumption that the agents directly head toward the target.
$v_{\max}^{2}-\|\dot{\textbf{x}}_{c}^{i}(t)\|^{2}\geq 0,\ a_{\max}^{2}-\|\ddot{\textbf{x}}_{c}^{i}(t)\|^{2}\geq 0,$ (23a)
$\dot{\psi}_{c}(t)=\frac{\det\big{(}[(\textbf{x}_{q}(t)-\textbf{x}_{c}^{i}(t))^{\top};(\dot{\textbf{x}}_{q}(t)-\dot{\textbf{x}}_{c}^{i}(t))^{\top}]\big{)}}{\|\textbf{x}_{q}(t)-\textbf{x}_{c}^{i}(t)\|^{2}},$ (23b)
$-\dot{\psi}_{\max}\leq\dot{\psi}_{c}(t)\leq\dot{\psi}_{\max}$ (23c)
@@ VI Validations
In this section, the proposed method is validated through various target-tracking settings. We measure the distance between the drone and environments, defined as (25a), the distance among agents (25b), and the distance between the target and drone, defined as (25c), to evaluate drone safety.
@@
$\chi_{3}(t)=\min_{i\in\mathcal{I}_{c}}\min_{\begin{subarray}{c}\textbf{x}(t)\in\mathcal{C}_{i}(t)\\ \textbf{y}(t)\in\mathcal{T}(t)\end{subarray}}\|\textbf{x}(t)-\textbf{y}(t)\|$ (25c)
Also, we measure the distance between the Lines-of-Sight and environments (26a) and the distance between the Lines-of-Sight and the other agents (26b) to assess the target visibility.
$\phi_{1}(t)=\min_{i\in\mathcal{I}_{c}}\min_{\begin{subarray}{c}\textbf{x}(t)\in\mathcal{L}(\textbf{x}_{c}^{i}(t),\textbf{x}_{q}(t))\\ \textbf{y}(t)\in\mathcal{O}(t)\end{subarray}}\|\textbf{x}(t)-\textbf{y}(t)\|,$ (26a)
$\phi_{2}(t)=\min_{i\in\mathcal{I}_{c}}\min_{\begin{subarray}{c}\textbf{x}(t)\in\mathcal{L}(\textbf{x}_{c}^{i}(t),\textbf{x}_{q}(t))\\ \textbf{y}(t)\in\mathcal{C}^{i}(t)\end{subarray}}\|\textbf{x}(t)-\textbf{y}(t)\|$ (26b)
Using the above performance metrics, we validate the operability of the proposed planner under two environmental conditions: 1) unstructured static obstacles, and 2) dynamic obstacles. Through simulations and hardware experiments, we show successful target tracking. Table II shows the reported performance, and the details of the tests are explained in the following subsections. In addition, we validate the effectiveness of the proposed method for multi-agent tracking through a comparative analysis and show applicability to 3D scenarios.
@@ VI-A Simulations
Scenario 1 (Unstructured Environment): We compare the proposed planner with the state-of-the-art planner [20] in an environment with various shapes of static obstacles, as shown in Fig. 4. As Table II shows, the trackers controlled by the baseline [20] experience inter-agent occlusions and collisions with the target several times, despite finely tuned parameters to adapt to the narrow tracking environment. In contrast, our planner successfully tracks the target without collision and occlusion by thoroughly checking safety and visibility constraints, including the tracking distance $(d_{\min},d_{\max})$ and DIVC constraints.
@@ VI-C Comparison Analysis
The performance of the proposed approach is investigated through two tests. The tests evaluate the effectiveness of the DBVC, the DIVC, and improved feasibility check methods. In each test, we define success as the absence of occlusion or collision until all objects come to a stop. We conduct 1000 tests for each tracking setup and measure the success rate.
First, we validate the effectiveness of the DBVC and the DIVC by comparing the success rates with cases where these two cells are not applied and where static versions of the cells are used. Specifically, when the cells are not used, each tracker treats the neighbor trackers as dynamic obstacles. In the case of using static cells, we use BVC, $\mathcal{H}_{\mu 1}^{ij}$, and $\mathcal{H}_{\mu 2}^{ij},\mu=o$ or $a$, instead of $\mathcal{H}_{s}^{ij}(t)$, $\mathcal{H}_{\mu 1}^{ij}(t)$, and $\mathcal{H}_{\mu 2}^{ij}(t)$.
@@
Second, we conduct a benchmark test to validate the superiority of our planner in a dynamic obstacle environment. We compare our method with a noncooperative approach and a conservative approach. For the noncooperative approach, we use the BPMP-Tracker [8]. Since the method in [8] is designed for single-agent tracking, we adapt the problem setting so that each tracker treats neighboring agents as obstacles. The conservative approach utilizes the DBVC and the DIVC but applies the conservative feasibility check methods from [8].
@@
Figure 6: Tracking in 3D spaces. Reported paths of trackers (blue), target (red), and dynamic obstacles (green), with static obstacles (black) and Line-of-sights (purple) connecting the trackers and the target.
VI-D 3D Extensions
The proposed planner, initially designed for 2D, is extended to 3D environments. Allowing trackers to move along the z-axis enables a wider range of feasible motions and improves tracking performance in challenging scenarios. The following tests show its effectiveness in 3D settings.
Scenario 5: (3D Unstructured Environment) The aerial target navigates though height-varied structures and tunnel-like passages while adjusting its altitude within the range of 1 to 2 meters, with the maximum speed of 2.19 m/s. The left side of Fig. 6 shows the paths of the target and the four trackers.
Scenario 6: (3D Dynamic Environment) The aerial target flies in a $6\times 6\times 4$ $\text{m}^{3}$ space with 100 moving airborne obstacles, with the maximum speed of $1.0$ m/s. The right side of Fig. 6 illustrates the results of tracking using four trackers.
@@ VII Conclusion
We presented a distributed multi-agent trajectory generation method for aerial tracking, which prevents both occlusion and collision caused by obstacles. DBVC and DIVC were constructed by dividing the space into multiple cells to avoid inter-agent collision and inter-agent occlusion, respectively. Since both cells are designed based on the shared current positions of agents and the predicted target’s trajectory, it enables distributed planning. We achieved fast computation by combining the DBVC and the DIVC with the Bernstein-polynomial-motion-primitive-based trajectory planner and validated the operability of our planner through tests under various tracking conditions. Lastly, we confirmed that the proposed method achieved a higher success rate in tracking missions in environments with dynamic obstacles, outperforming the state-of-the-art methods.